# Lindbladian approach to $P(E)$ theory

### v1

@@ V Electric current
In this section, we calculate the current flowing through the junction based on the Lindblad equation (12). Since the Lindbladian contains only particle operators, the current must also be expressed in terms of particle operators rather than circuit variables. We start with the expectation value of
@@
is the operator for the total charge in the left lead. The expectation value of the current flowing from left to right is calculated as
@@
We substitute the right-hand side of the Lindblad equation into $\partial_{t}\rho_{S}(t)$, make use of the cyclic properties of the trace and the relation $\left[H_{0}+H_{LS},c_{Lk\sigma}^{+}c_{Lk\sigma}\right]=0$. We obtain
@@
In this section, we presented the relation between the conventional approach to the $P(E)$ theory and the Lindbladian formalism. We have seen that the assumptions of Fermi’s golden rule are essentially equivalent to the Born-Markovian approximation used for the Lindblad equation. Namely, both approaches are valid if the coupling between the system and the bath is weak and the results are given up to leading order only. Furthermore, both approaches assume that the tunneling events are independent from each other and that the reservoir completely relaxes between consecutive events.
VI Dissipated heat
@@
The dissipated heat is a fundamentally important quantity [33]. Furthermore, it can be used to characterize devices through Joule spectrometry, e.g. in hybrid superconducting devices [34]. In this section, we study the heat current flowing through the junction with normal-state leads. The total energy of the left lead is given by the operator $H_{L}=\sum_{k\sigma}\varepsilon_{k}c_{Lk\sigma}^{+}c_{Lk\sigma}$ and its expectation value is given by
@@
where we used again that the leads are in thermal equilibrium. Similar formula can be derived for the heat flow to the right lead.
@@
Without environmental effects, $P(E)=\delta(E)$, the two heat currents are the same, $P_{L}=P_{R}$. In the presence of the photonic bath, however, some part of the heat current which leaves the left lead does not arrive on the right lead but gets dissipated into the environment. The dissipated heat is calculated as
@@
In the formula, $R_{T}$ is the electric resistance of the junction as given in Eq. (23). The dissipated heat has been computed numerically for an $RLC$ environment, see Fig. 2. The circuits are distinguished based on the quality factor defined as $Q_{f}=\sqrt{L/C}/R$ following Ref. [20]. In each case, $L=4C\hbar^{2}/e^{2}$ has also been set. The quality factor indicates how pronounced are the resonant peaks in the $P(E)$ function. For lower values of $Q_{f}$, the ohmic resistance becomes more dominant.
Figure 2: Derivative of the dissipated power as a function of bias voltage for three different values of the quality factor $Q_{f}=\sqrt{L/C}/R$ (numerical results). The inset shows the corresponding $P(E)$ functions in units of $E_{C}^{-1}$ where $E_{C}=e^{2}/(2C)$ is the charging energy. The temperature is $T=0.01E_{C}/k_{\mathrm{B}}$.
At low bias voltage, the dissipated heat has a rich structure and depends strongly on the specific circuit. At high voltages, however, the dissipated heat behaves as $P_{diss}\sim V$ which is noticeably distinct from the quadratic behavior characteristic of ohmic resistances. This result is particularly interesting since the $I(V)$ characteristics do exhibit the ohmic behavior in this regime, $I=V/R_{T}$. It can also be shown that in the large bias voltage limit, the dissipated heat through the junction is given by
@@
We remark that the dissipated heat is not necessarily lost on the ohmic resistance of the electric circuit. It can be shown that even in the case of a simple $LC$ circuit, the dissipated heat is non-zero. This can be explained by that the photonic environment with a well-defined temperature is coupled to a large heat bath and the dissipated heat can also be lost through this coupling.
VII Conclusion
The $P(E)$ theory is the most commonly used approach to explain the transport properties of ultrasmall junctions. In this paper we have revealed the relation between the conventional derivation based on Fermi’s golden rule and the Lindbladian formalism. We emphasize that the Lindbladian approach leads to the same result on the level of $I(V)$ characteristics. The Lindbladian approach, however, enables one to include in consideration further effects such as additional dissipative features or tunneling processes without interaction with the photons. Furthermore, the Lindbladian approach also enables the application of the multitude of numerical techniques designed for solving the Lindblad equations to the tunneling phenomena.
One of the main results of the present paper is the demonstration that the infinite-time limit of Fermi’s golden rule and the assumption of the independence of the tunneling events are essentially equivalent to the Markovian approximation applied when deriving the Lindblad equation. We believe that our results extend the conventional $P(E)$ theory, open the route to more complex systems, and enable the calculation of more complicated observables. As a demonstration, we have applied the formalism to compute the heat currents through the junction. The dissipated heat exhibits a $P_{\mathrm{diss}}\sim V$ behaviour at large bias voltage even though the current obeys $I=V/R_{T}$ in this regime.
Acknowledgements.
We acknowledge the support of the Slovenian Research and Innovation Agency (ARIS) under P1-0416 and J1-3008. Á.B. acknowledges the support of the National Research, Development and Innovation Office - NKFIH Project No. K142179.

### v2

@@ V Electric current
In this section, we calculate the current flowing through the junction based on the Lindblad equation (21). Since the Lindbladian contains only particle operators, the current must also be expressed in terms of particle operators rather than circuit variables. We start with the expectation value
@@
is the operator for the total charge in the left lead. The expectation value of the net current flowing from left to right is calculated as
@@
We substitute the right-hand side of the Lindblad equation into $\partial_{t}\rho_{S}(t)$, make use of the cyclic properties of the trace and the relation $\left[H_{0}+H_{LS},c_{Lk\sigma}^{+}c_{Lk\sigma}\right]=0$ which leads to
@@
Based on the Lindblad equation, the heat currents can also be computed by taking the time-derivative of the expectation value of the total energy in the left (right) lead $H_{L}$ ($H_{R}$), for details see Appendix A. Due to the energy exchange between the electrons and the photonic field, the heat current leaving the left lead $P_{L}$ differs from the heat current arriving to the right lead, $P_{R}$. The difference is the dissipated heat, $P_{\mathrm{diss}}=P_{L}-P_{R}$, which is obtained from the Lindblad equation as
$\begin{split}&P_{\mathrm{diss}}=\frac{2\pi}{\hbar}\sum_{kq\sigma}(\varepsilon_{k}-\varepsilon_{q})|T_{kq}|^{2}\times\\ &\times\Big{(}P(\varepsilon_{k}-\varepsilon_{q})f_{L}(1-f_{R})-P(\varepsilon_{q}-\varepsilon_{k})(1-f_{L})f_{R}\Big{)}.\end{split}$ (33)
which is in accordance with Refs. [43, 44, 45]. This further confirms the validity of the Lindbladian formalism based on Eq. (21).
@@
In this section, we presented the relation between the conventional approach to the $P(E)$ theory and the Lindbladian formalism based on tracing out the photonic degrees of freedom. We have seen that the assumptions of Fermi’s golden rule are essentially the same as the Born-Markovian approximation used for the Lindblad equation. Namely, both approaches are valid if the coupling between the system and the bath is weak and the results are given up to leading order only. Furthermore, both approaches assume that the tunneling events are independent from each other and that the reservoir completely relaxes between consecutive events.
VI Inelastic quasiparticle tunneling through Josephson junctions
We apply the Lindbladian formalism to the quasiparticle tunneling between superconducting leads. Note that this phenomenon is distinct from the tunneling of Cooper pairs [21, 46] which is not covered in this section. In Ref. 21, the quasiparticle tunneling is handled such that the only difference to the normal leads comes from the difference in the density of states of elementary excitations. Ref. 24, however, claims that due to the fact that superconducting elementary excitations are linear superpositions of particles and holes with suitable energy-dependent coefficients, the corresponding coherence factors appear in the expressions for the tunneling rates. Here, we show that by taking into account the Cooper pair counting in the Bogoliubov operators, the results of Ref. 21 are recovered in special cases but coherence factors appear in more general situations.
Let us start with the Hamiltonian of the superconducting (left) lead which is obtained within mean-field approximation as
$H_{L}-\mu_{L}\hat{N}_{L}=\sum_{k\sigma}E_{L}(k)d_{Lk\sigma}^{+}d_{Lk\sigma}$ (34)
where $\hat{N}_{L}$ is the total number of particles, $E_{L}(k)=\sqrt{\xi_{L}(k)^{2}+|\Delta_{L}|^{2}}$ is the energy spectrum with $\Delta_{L}$ the superconducting gap and $\xi_{L}(k)=\varepsilon_{L}(k)-\mu_{L}$. In the formula, $d_{Lk\sigma}$ is the annihilation operator of the Bogoliubov quasi-particles. Following Refs. [47, 48, 49], the annihilation operators are defined as
$d_{Lk\sigma}=u_{k}c_{Lk\uparrow}-\sigma v_{k}S_{L}c_{L,-k\bar{\sigma}}^{+},$ (35)
where we use the convention that the value of $\sigma$ is +1 for spin up and -1 for spin down and $\bar{\sigma}$ denotes the opposite of spin $\sigma$. Furthermore,
$u_{k}=\frac{1}{\sqrt{2}}\sqrt{1+\frac{\xi_{L}(k)}{E_{L}(k)}}\qquad v_{k}=\frac{e^{i\phi}}{\sqrt{2}}\sqrt{1-\frac{\xi_{L}(k)}{E_{L}(k)}}$ (36)
where $\phi$ is the phase of $\Delta_{L}$. The non-standard feature of Eq. (35) is the inclusion of the operator $S_{L}$ which annihilates a Cooper-pair from the the left lead. Technically, we consider that $S_{L}$ acts on an auxiliary Hilbert space of Cooper-pairs. The basis elements of this Hilbert space are $|M_{L}\rangle$ describing a state with $M_{L}$ Cooper pairs and fulfilling $S_{L}|M_{L}\rangle=|M_{L}-1\rangle$. The overall state of the superconducting lead takes the form of a tensor product, $|\mathrm{electrons}\rangle\otimes|M_{L}\rangle$.
As pointed out in Refs. [47, 48, 49], the application of $S_{L}$ ensures proper accounting for the electric charge, which is necessary for the discussion of tunneling problems. Indeed, as shown in Appendix B, the annihilation operators fulfill the relation
$\left[\hat{N}_{L},d_{Lk\sigma}\right]=-d_{Lk\sigma}$ (37)
if the total number of particles operator is defined as
$\hat{N}_{L}=\sum_{k\sigma}c_{Lk\sigma}^{+}c_{Lk\sigma}+2\sum_{M_{L}}M_{L}|M_{L}\rangle\langle M_{L}|,$ (38)
where the factor 2 stems from the fact that Cooper pairs carry a charge of $2e$.
For the right lead, similar expressions can be derived and with the Hamiltonian $H_{R}$, the total Hamiltonian of the leads is given as $H_{0}=H_{L}+H_{R}$. Using the Bogoliubov transformation (35), the tunneling Hamiltonian of (3) is rewritten as
$H_{T}=\sum_{kq\sigma}\Big{[}T_{kq}e^{-i\varphi}\left(u_{k}u_{q}d_{q\sigma}^{+}d_{k\sigma}-v_{q}^{*}v_{k}S_{R}^{+}S_{L}d_{k\sigma}^{+}d_{q\sigma}+\sigma v_{q}^{*}u_{k}S_{R}^{+}d_{q\bar{\sigma}}d_{k\sigma}+\sigma u_{q}v_{k}S_{L}d_{q\sigma}^{+}d_{k\bar{\sigma}}^{+}\right)+$
$+T_{kq}^{*}e^{i\varphi}\left(u_{k}u_{q}d_{k\sigma}^{+}d_{q\sigma}-v_{q}v_{k}^{*}S_{L}^{+}S_{R}d_{q\sigma}^{+}d_{k\sigma}+\sigma v_{k}^{*}u_{q}S_{L}^{+}d_{k\bar{\sigma}}d_{q\sigma}+\sigma u_{k}v_{q}S_{R}d_{k\sigma}^{+}d_{q\bar{\sigma}}^{+}\right)\Big{]},$ (39)
from which the jump processes are identified as shown in Table 2 in accordance with Ref. 48. In the table, the frequencies are determined based on the energy difference between the initial and the final states of the process. For the energy, we consider the eigenvalues of $H_{L}$ and $H_{R}$ and not $H_{L}-\mu_{L}\hat{N}_{L}$ or $H_{R}-\mu_{R}\hat{N}_{R}$. That is, we express $H_{L}$ from Eq. (34) and obtain $H_{R}$ similarly. This procedure gives rise to the terms $\mu_{L}-\mu_{R}=eV$ in the frequencies of the jump processes.
id $A_{p}$ description frequency
1+ $T_{kq}u_{k}u_{q}d_{q\sigma}^{+}d_{k\sigma}$ $L\rightarrow R$, C: - $\hbar\omega_{p}=E_{k}+eV-E_{q}$
2+ $-T_{kq}v_{q}^{*}v_{k}S_{R}^{+}S_{L}d_{k\sigma}^{+}d_{q\sigma}$ $R\rightarrow L$, C: $L\rightarrow R$ $\hbar\omega_{p}=E_{q}+eV-E_{k}$
3+ $T_{kq}v_{q}^{*}u_{k}S_{R}^{+}d_{q\bar{\sigma}}d_{k\sigma}$ $LR\rightarrow\varnothing$, C: $\varnothing\rightarrow R$ $\hbar\omega_{p}=E_{k}+E_{q}+eV$
4+ $T_{kq}u_{q}v_{k}S_{L}d_{q\sigma}^{+}d_{k\bar{\sigma}}^{+}$ $\varnothing\rightarrow LR$, C: $L\rightarrow\varnothing$ $\hbar\omega_{p}=eV-E_{k}-E_{q}$
1- $T_{kq}^{*}u_{k}u_{q}d_{k\sigma}^{+}d_{q\sigma}$ $R\rightarrow L$, C: - $\hbar\omega_{p}=E_{q}-eV-E_{k}$
2- $-T_{kq}^{*}v_{q}v_{k}^{*}S_{L}^{+}S_{R}d_{q\sigma}^{+}d_{k\sigma}$ $L\rightarrow R$, C: $R\rightarrow L$ $\hbar\omega_{p}=E_{k}-eV-E_{q}$
3- $T_{kq}^{*}u_{k}v_{q}S_{R}d_{k\sigma}^{+}d_{q\bar{\sigma}}^{+}$ $\varnothing\rightarrow LR$, C: $R\rightarrow\varnothing$ $\hbar\omega_{p}=-eV-E_{k}-E_{q}$
4- $T_{kq}^{*}v_{k}^{*}u_{q}S_{L}^{+}d_{k\bar{\sigma}}d_{q\sigma}$ $LR\rightarrow\varnothing$, C: $\varnothing\rightarrow L$ $\hbar\omega_{p}=E_{k}+E_{q}-eV$
Table 2: Jump processes in the case of superconducting leads.
It is worth to notice the critical importance of the $S_{L,R}$ operators. For example, the processes 1+ and 2- both describe a quasiparticle tunneling from left to right. However, in the process 2-, it is accompanied by a hopping of a Cooper pair from right to left producing a net charge transfer from right to left.
To calculate the electric current through the junction, we follow the same procedure as presented in Sec. V. The total charge of the left lead is given by $\hat{Q}_{L}=e\hat{N}_{L}$ with the total number of particles defined in Eq. (38). Formally, we obtain the same result as Eq. (27), but the sum over the different processes includes now all processes of Table 2. We assume again that the density matrix $\rho_{S}(t\rightarrow\infty)$ describes a thermal equilibrium state of the superconducting leads. As a result, only the diagonal terms $p=p^{\prime}$ contribute to the sum in Eq. (27) leading to
$I=\frac{2\pi e}{\hbar}\sum_{kq\sigma}|T_{kq}|^{2}\Big{[}u_{k}^{2}u_{q}^{2}f_{L}(1-f_{R})P(E_{L}(k)-E_{R}(q)+eV)+|v_{k}|^{2}|v_{q}|^{2}f_{R}(1-f_{L})P(E_{R}(q)-E_{L}(k)+eV)+$
$+u_{k}^{2}|v_{q}|^{2}f_{L}f_{R}P(E_{L}(k)+E_{R}(q)+eV)+u_{q}^{2}|v_{k}|^{2}(1-f_{L})(1-f_{R})P(-E_{L}(k)-E_{R}(q)+eV)-$
$-u_{k}^{2}u_{q}^{2}(1-f_{L})f_{R}P(E_{R}(q)-eV-E_{L}(k))-|v_{k}|^{2}|v_{q}|^{2}f_{L}(1-f_{R})P(E_{L}(k)-eV-E_{R}(q))-$
$-|v_{q}|^{2}u_{k}^{2}(1-f_{R})(1-f_{L})P(-eV-E_{L}(k)-E_{R}(q))-u_{q}^{2}|v_{k}|^{2}f_{L}f_{R}P(E_{L}(k)+E_{R}(q)-eV)\Big{]}$ (40)
where $f_{L}=f(E_{L}(k))$ and $f_{R}=f(E_{R}(q))$.
In the case when both $|T_{kq}|^{2}$ and the density of states of $\xi_{L/R}$ are independent of the wavenumber and energy, we obtain
$I=\frac{1}{eR_{T}}\int_{-\infty}^{\infty}\mathrm{d}\xi_{L}\int_{-\infty}^{\infty}\mathrm{d}\xi_{R}\Big{[}u_{k}^{2}u_{q}^{2}f_{L}(1-f_{R})P(E_{L}-E_{R}+eV)+|v_{k}|^{2}|v_{q}|^{2}f_{R}(1-f_{L})P(E_{R}-E_{L}+eV)+$
$+u_{k}^{2}|v_{q}|^{2}f_{L}f_{R}P(E_{L}+E_{R}+eV)+u_{q}^{2}|v_{k}|^{2}(1-f_{L})(1-f_{R})P(-E_{L}-E_{R}+eV)-$
$-u_{k}^{2}u_{q}^{2}(1-f_{L})f_{R}P(E_{R}-eV-E_{L})-|v_{k}|^{2}|v_{q}|^{2}f_{L}(1-f_{R})P(E_{L}-eV-E_{R})-$
$-|v_{q}|^{2}u_{k}^{2}(1-f_{R})(1-f_{L})P(-eV-E_{L}-E_{R})-u_{q}^{2}|v_{k}|^{2}f_{L}f_{R}P(E_{L}+E_{R}-eV)\Big{]},$ (41)
where $E_{L/R}=\sqrt{\xi_{L/R}^{2}+|\Delta_{L/R}|^{2}}$ and $R_{T}$ is the resistance given in Eq.  (32). One can observe that only the coherence factors, e.g., $u_{k}^{2}|v_{q}|^{2}$, depend on the sign of $\xi_{L/R}$. We analyze all four sectors of the $(\xi_{L},\xi_{R})$ plane and sum up with respect to the sign of $\xi_{L}$ and $\xi_{R}$. By using the functions $n_{L/R}(E)=\Theta(|E|-|\Delta_{L/R}|)|E|/\sqrt{E^{2}-|\Delta_{L/R}|^{2}}$ with the Heaviside-function $\Theta(x)$, we obtain
$I=\frac{1}{eR_{T}}\int_{-\infty}^{\infty}\mathrm{d}E\int_{-\infty}^{\infty}\mathrm{d}E^{\prime}n_{L}(E)n_{R}(E^{\prime})\times$
×$\times\Big{[}f(E)(1-f(E^{\prime}))P(E-E^{\prime}+eV)-$
$f(E^{\prime})(1-f(E))P(E^{\prime}-E-eV)\Big{]},$ (42)
which is the same as Eq. (153) in Ref. 21. Note that no coherence factors appear in this expression.
Let us now study the case of non-constant density of states and assume linear energy dependence in the vicinity of the chemical potential:
$\begin{split}\mathcal{D}_{L}(\xi_{k})&=\mathcal{D}_{L0}\left(1+s_{L}\xi_{k}\right),\\ \mathcal{D}_{R}(\xi_{q})&=\mathcal{D}_{R0}\left(1+s_{R}\xi_{q}\right).\end{split}$ (43)
Note that such energy dependence of $\mathcal{D}$ implies that the particle-hole symmetry is broken in the superconducting leads.
By substituting the density of states into (40) and following the same procedure as previously presented for the constant density case, we now obtain
$I=\frac{1}{eR_{T}}\int_{-\infty}^{\infty}\mathrm{d}E\int_{-\infty}^{\infty}\mathrm{d}E^{\prime}n_{L}(E)n_{R}(E^{\prime})\times$
×$\times\left(1+s_{L}\frac{E^{2}-|\Delta_{L}|^{2}}{E}\right)\left(1+s_{R}\frac{E^{\prime 2}-|\Delta_{R}|^{2}}{E^{\prime}}\right)\times$
×$\times\Big{[}f(E)(1-f(E^{\prime}))P(E-E^{\prime}+eV)-$
$f(E^{\prime})(1-f(E))P(E^{\prime}-E-eV)\Big{]},$ (44)
where $R_{T}$ is defined as Eq. (32) but with $\mathcal{D}_{L/R0}$.
The integrals of Eq. (44) are evaluated numerically and the results are shown with solid lines in Fig. 2 for an $RLC$ circuit environment where $Q_{f}=\sqrt{L/C}/R=0.25$ and $L=4C\hbar^{2}/e^{2}$ following Sec. 4.3 in Ref. 21. The numerical calculations have been performed in two distinct cases. In the first (parity symmetric) case, the density of states behave the same in both leads, $s_{L}=s_{R}=s$, while in the second (parity anti-symmetric) case, the slopes of the energy dependencies have opposite sign, $s_{L}=-s_{R}=s$. The parity anti-symmetric case exhibits more significant $s$-dependence.
At higher bias voltages, the $I(V)$ characteristics converges to the curve corresponding to the case of no environment and no superconductivity. For $s_{L,R}=0$, this reference $I(V)$ function is a linear describing Ohmic behavior. For non-zero $s_{L,R}$, the reference current is calculated as
$\begin{split}I_{0}(V)&=\frac{V}{R_{T}}\Bigl{(}1+(s_{R}-s_{L})\frac{eV}{2}\\ &-\frac{s_{L}s_{R}}{6}\left((eV)^{2}-2\pi^{2}(k_{B}T)^{2}\right)\Bigr{)}\end{split}$ (45)
which is plotted with dashed lines in Fig. 2. Note that the deviation of the $I(V)$ curves from the Ohmic behaviour at high voltages is not a consequence of the inelastic tunneling but follows from the broken particle-hole symmetry.
The result of Eq. (44) demonstrates how the broken particle-hole symmetry occurs in the $I(V)$ characteristics. The coherence factors $\left(1+s\frac{E^{2}-|\Delta|^{2}}{E}\right)$ could be derived only by properly identifying the charge conserving jump processes within the Lindbladian formalism.
Figure 2: $I(V)$ characteristics for different types of density of states. The case of uniform density of states (blue, solid) is compared to the parity symmetric case, $s_{L}=s_{R}=s_{\mathrm{sym}}=0.05$ (orange, solid), and to the parity anti-symmetric case, $s_{L}=s_{R}=s_{\mathrm{asym}}=0.05$ (green, solid). The dashed lines represent Eq. (45).
VII Conclusion
The $P(E)$ theory is the most commonly used approach to explain the transport properties of ultrasmall tunnel junctions. In this paper, we showed that the conventional approach to $P(E)$ theory, which is based on Fermi’s golden rule, relies on essentially the same assumptions as the Born-Markovian approximation of the Lindblad equation approach. The Lindbladian formalism presented here involves tracing out only photonic degrees of freedom. This has to be contrasted with the typical approaches in the literature where the leads are treated as reservoirs. The main advantage of keeping all electronic degrees of freedom as system variables is that further effects can be included in the model, such as additional dissipative processes or elastic tunneling terms. Furthermore, additional scattering processes could also be taken into consideration. A particularly interesting situation could be to consider an Anderson impurity at the edge of one or both leads and study the interplay between the dissipation and the Kondo effect.
We believe that our results extend the conventional $P(E)$ theory, open the route to more complex systems, and enable the calculation of more complicated observables. As a demonstration, we have applied the formalism to inelastic quasiparticle tunneling through superconducting junctions. It has been shown that if the normal-state density of states of the leads is flat (constant in energy), no coherence factors occur in the $I(V)$ characteristics. For a non-constant density of states (i.e., broken particle-hole symmetry), however, the coherence factors play an important role.
Acknowledgements.
We thank the anonymous referee for their valuable insights. We acknowledge the support of the Slovenian Research and Innovation Agency (ARIS) under P1-0416 and J1-3008. Á.B. acknowledges the support of the National Research, Development and Innovation Office - NKFIH Project No. K142179.
Appendix A Dissipated heat
@@
The dissipated heat is a fundamentally important quantity [50]. Furthermore, it can be used to characterize devices through Joule spectrometry, e.g. in hybrid superconducting devices [51]. In this section, we study the heat current flowing through the junction with normal-state leads. The heat transport has been broadly studied in the literature [43, 44, 45]. Here, we demonstrate how these results can be recovered within the Lindbladian formalism.
The total energy of the left lead is given by the operator $H_{L}=\sum_{k\sigma}\varepsilon_{k}c_{Lk\sigma}^{+}c_{Lk\sigma}$ and its expectation value is given by
@@
where we again took into account that the leads are in thermal equilibrium. Similar formula can be derived for the heat flow to the right lead:
@@
In the absence of environmental effects, when $P(E)=\delta(E)$, the two heat currents are the same, $P_{L}=P_{R}$. In the presence of the photonic bath, however, some part of the heat current which leaves the left lead does not arrive on the right lead but gets dissipated into the environment. The dissipated heat is calculated as
@@
In the formula, $R_{T}$ is the electric resistance of the junction as given in Eq. (32).
The dissipated heat has been computed numerically for an $RLC$ environment, see Fig. 3. The circuits are distinguished based on the quality factor defined as $Q_{f}=\sqrt{L/C}/R$ following Ref. [21]. In each case, $L=4C\hbar^{2}/e^{2}$ has also been set. The quality factor indicates how pronounced are the resonant peaks in the $P(E)$ function. For lower values of $Q_{f}$, the ohmic resistance becomes dominant.
Figure 3: Derivative of the dissipated power as a function of bias voltage for three different values of the quality factor $Q_{f}=\sqrt{L/C}/R$ (numerical results). The inset shows the corresponding $P(E)$ functions in units of $E_{C}^{-1}$ where $E_{C}=e^{2}/(2C)$ is the charging energy. The temperature is $T=0.01E_{C}/k_{\mathrm{B}}$. The $P(E)$ curves are the same as in Sec. 4.3 of Ref. [21].
At low bias voltage, the dissipated heat has a rich structure and depends strongly on the specific circuit. At high voltages, however, the dissipated heat behaves as $P_{\mathrm{diss}}\sim V$ which is noticeably distinct from the quadratic behavior characteristic of ohmic resistances. This result is particularly interesting since the $I(V)$ characteristics do exhibit the ohmic behavior in this regime, $I=V/R_{T}$. It can also be shown that in the large bias voltage limit, the dissipated heat through the junction is given by
@@
We remark that the dissipated heat is not necessarily lost on the ohmic resistance of the electric circuit. It can be shown that even in the case of a simple $LC$ circuit, the dissipated heat is non-zero. This can be explained by the fact that the photonic environment with a well-defined temperature is, by definition, coupled to a large heat bath and the dissipated heat can also be lost through this coupling.
Appendix B Tracking the number of particles in superconducting leads
When studying tunneling phenomena [49, 48, 47], it is beneficial to introduce the Coulomb pair operators $S$ in the annihilation operators of Bogoliubov quasiparticles. These operators track the number of Cooper pairs forming the superconducting condensate: $S$ decreases that number by one, $S^{+}$ increases it by one. The annihilation operators of the Bogoliubov quasiparticles are then given as
$d_{k\sigma}=u_{k}c_{k\sigma}-\sigma v_{k}Sc_{-k\bar{\sigma}}^{+}$ (53)
where we use the convention that the value of $\sigma$ is +1 for spin up and -1 for spin down and $\bar{\sigma}$ denotes the opposite of spin $\sigma$. The coefficients are given by the standard expressions
$\begin{split}u_{k}&=\frac{1}{\sqrt{2}}\sqrt{1+\frac{\xi_{k}}{E_{k}}},\\ v_{k}&=\frac{e^{i\varphi}}{\sqrt{2}}\sqrt{1-\frac{\xi_{k}}{E_{k}}},\end{split}$ (54)
with $\xi_{k}=\varepsilon(k)-\mu$ and $E_{k}=\sqrt{\xi_{k}^{2}+|\Delta|^{2}}$. The operator $S$ annihilates a Cooper pair from the system, hence, under the action of $d_{k\sigma}$ the charge of the total system decreases by exactly one unit. In this section, we consider only one lead and we have hence dropped the lead index.
The inverse Bogoliubov transformation is given by
$c_{k\sigma}=u_{k}d_{k\sigma}+\sigma v_{k}Sd_{-k\bar{\sigma}}^{+},$ (55)
which is valid if $SS^{+}=1$. This relation will be checked later. An interesting property of Eq. (53) is that it does not diagonalize the BCS mean-field Hamiltonian in the standard form
$\begin{split}H_{\mathrm{BCS}}&=\sum_{k}\xi_{k}\left(c_{k\uparrow}^{+}c_{k\uparrow}+c_{-k\downarrow}^{+}c_{-k\downarrow}\right)\\ &-\sum_{k}\left(\Delta c_{k\uparrow}^{+}c_{-k\downarrow}^{+}+h.c.\right)\end{split}$ (56)
because the terms of the type $c^{+}c^{+}$, which change the total charge by $2e$, cannot be expressed in terms of $d^{+}d$, which do not change the total charge. Instead, one has to revisit the mean-field approximation applied to the original Hamiltonian
$\begin{split}H&=\sum_{k}\xi_{k}\left(c_{k\uparrow}^{+}c_{k\uparrow}+c_{-k\downarrow}^{+}c_{-k\downarrow}\right)\\ &-\frac{g}{N}\sum_{kk^{\prime}}c_{k\uparrow}^{+}c_{-k\downarrow}^{+}c_{-k^{\prime}\downarrow}c_{k^{\prime}\uparrow}.\end{split}$ (57)
Let us insert an identity operator $SS^{+}$ as
$\begin{split}H&=\sum_{k}\xi_{k}\left(c_{k\uparrow}^{+}c_{k\uparrow}+c_{-k\downarrow}^{+}c_{-k\downarrow}\right)\\ &-\frac{g}{N}\sum_{kk^{\prime}}c_{k\uparrow}^{+}c_{-k\downarrow}^{+}SS^{+}c_{-k^{\prime}\downarrow}c_{k^{\prime}\uparrow}\end{split}$ (58)
and assume that the following expectation value
$\Delta=\frac{g}{N}\sum_{k}\langle S^{+}c_{-k\downarrow}c_{k\uparrow}\rangle$ (59)
is non-zero in the ordered phase. By performing the standard steps of the mean-field approximation, we obtain
$H_{MF}-\frac{N|\Delta|^{2}}{g}=$
$\sum_{k}\xi_{k}\left(c_{k\uparrow}^{+}c_{k\uparrow}+c_{-k\downarrow}^{+}c_{-k\downarrow}\right)-\sum_{k}\left(\Delta c_{k\uparrow}^{+}c_{-k\downarrow}^{+}S+h.c.\right)$ (60)
which is readily diagonalized by the transformation (53) resulting in Eq. (34). It is remarkable that, in contrast to (56), the Hamiltonian of (60) conserves the total charge because the terms such as $c^{+}c^{+}S$ describe the creation of two electrons with the simultaneous annihilation of a Cooper pair. Similar description has been recently used in Ref. 46.
Technically, by this procedure we have introduced an auxiliary Hilbert space for the Cooper pairs which is separate from the physical electron Hilbert space. This redundancy in description serves to track the number of electrons, which is important in discussing tunneling between multiple superconductors. The states of the system take the following general form:
$|\Psi\rangle=|\mbox{\{many body state of electrons\}}\rangle\otimes|M\rangle_{c}$ (61)
where the last part, $|M\rangle_{c}$, describes a state with $M$ Cooper-pairs. In the auxiliary space, the basis vectors are $|M\rangle_{c}$ with integer $M$ ranging from 0 to infinity.
The $S$ operator acts as
$S=\mathrm{I}\otimes\sum_{M=0}^{\infty}|M\rangle_{c}\langle M+1|_{c}$ (62)
from which $SS^{+}=1$ also follows.
The ground state of the Hamiltonian in Eq. (60) is
$|\mathrm{GS}\rangle=\prod_{k}\left(u_{k}+v_{k}Sc_{k\uparrow}^{+}c_{-k\downarrow}^{+}\right)|0\rangle\otimes|M_{0}\rangle_{c},$ (63)
where $|0\rangle$ is the vacuum (empty) state of electrons, while $M_{0}$ is half the total number of electrons in the superconductor. Note that the number of particles is the same in each term of the ground state, $N_{e}=2M_{0}$. Based on the ground state, the gap equation is derived as
$\Delta=\frac{g}{2N}\sum_{k}\frac{\Delta}{E_{k}}$ (64)
which coincides with the one in the standard BCS theory [49]. Finally, we express the operator for the total number of particles by considering both the electron and the Cooper pair contributions as
$\hat{N}=\sum_{k\sigma}c_{k\sigma}^{+}c_{k\sigma}+2\sum_{M}M|M\rangle\langle M|,$ (65)
which satisfies both $[\hat{N},S]=-2S$ and $[\hat{N},d_{k\sigma}]=-d_{k\sigma}$ in accordance with the physical intuition. Note that these commutators are crucial to properly track the charges which is essential when studying transport properties of superconducting junctions.